# https://github.com/yfzhang114/LLaVA-Align Debiasing Large Visual Language Models

### v1

@@ 3.3 Other Findings and Analysis
Navigating Challenges in Debiasing LLM-Biased Benchmarks: As illustrated in Fig. 10, we unravel the intricate interplay between debiasing methods and benchmarks influenced by LLM biases, with Naive’s performance centered around textual information. The post-hoc debiasing methods proposed exhibit inconsistent enhancements on the MMMU benchmark. Upon meticulous examination, when we entirely remove vision information by replacing the input image with an empty string (None) or the <unk> identifier (Unk), the model maintains comparable or superior performance in specific subsets. In these scenarios, our post-hoc methods, crafted to alleviate biases in LVLMs, fall short of yielding optimal results. Especially in instances where pure LLMs outshine, the application of post-hoc methods may prove detrimental. It is crucial to emphasize that benchmarks tailored for LVLMs should prioritize reliance on input images rather than solely on textual content. Consequently, our methods can act as valuable indicators; when the proposed debiasing methods exhibit suboptimal performance, it signals that the benchmark may tilt more towards LLMs and may not be apt for evaluating LVLMs effectively. In such cases, the proposed sampling methods, VDD, also demonstrate inferiority compared to Naive, with an average accuracy of versus .
@@ 0.C.1 Qualitative examples
We present examples in Fig. 15 to illustrate that the default decoding strategy leads to significant hallucinations in some instances, resulting in a low GPT-4 evaluation score (2). In these cases, VDD proves highly effective in alleviating hallucinations and achieving higher scores. A similar pattern is observed in Fig. 16, indicating that merely scaling up the model cannot completely eliminate unrealistic patterns, while VDD consistently delivers superior performance. Examples in Fig. 17 further affirm the efficacy and accuracy of VDD, generating more precise illustrations.
@@ Appendix 0.D Limitations and Future Work
| LLaVA-v1.5-7B | |||||||||
| Subsets | Naive | Post-Hoc Debias | Debias Sampling | ||||||
| None | Unk | Both | VCD[21] | VDD-None | VDD-UNK | Both | |||
| Object-level Hallucination | existence | 175.00 | 190.00 | 190.00 | 183.33 | 185.00 | 185.00 | 185.00 | 190.00 |
| count | 138.33 | 145.00 | 138.33 | 120.00 | 118.33 | 131.67 | 133.33 | 138.33 | |
| Attribute-level Hallucination | position | 128.33 | 145.00 | 140.00 | 130.00 | 128.33 | 131.67 | 133.33 | 126.67 |
| color | 131.67 | 170.00 | 165.00 | 145.00 | 155.00 | 165.00 | 158.33 | 165.00 | |
| Reasoning | commonsense | 102.86 | 95.00 | 94.29 | 105.00 | 102.14 | 92.14 | 120.00 | 113.57 |
| numerical | 60.00 | 70.00 | 67.50 | 45.00 | 92.50 | 70.00 | 77.50 | 70.00 | |
| translation | 92.50 | 50.00 | 50.00 | 47.50 | 70.00 | 72.50 | 92.50 | 75.00 | |
| code | 62.50 | 87.50 | 87.50 | 85.00 | 47.50 | 45.00 | 47.50 | 87.50 | |
| MME-OH | 313.33 | 335.00 | 328.33 | 303.33 | 303.33 | 316.67 | 318.33 | 328.33 | |
| MME-AH | 260.00 | 315.00 | 305.00 | 275.00 | 283.33 | 296.67 | 291.67 | 291.67 | |
| Reasoning | 317.86 | 302.50 | 299.29 | 282.50 | 312.14 | 279.64 | 337.50 | 346.07 | |
| LLaVA-v1.5-13B | |||||||||
| Subsets | Naive | Post-Hoc Debias | Debias Sampling | ||||||
| None | Unk | Both | VCD[21] | VDD-None | VDD-UNK | Both | |||
| Object-level Hallucination | existence | 180.00 | 188.33 | 188.33 | 190.00 | 173.33 | 190.00 | 188.33 | 190.00 |
| 111.67 | 155.00 | 155.00 | 145.00 | 143.33 | 135.00 | 163.33 | 133.33 | ||
| Attribute-level Hallucination | position | 130.00 | 165.00 | 165.00 | 155.00 | 123.33 | 138.33 | 143.33 | 143.33 |
| 150.00 | 155.00 | 155.00 | 130.00 | 130.00 | 160.00 | 165.00 | 165.00 | ||
| Reasoning | commonsense | 112.14 | 115.00 | 112.86 | 122.14 | 134.29 | 115.71 | 116.43 | 111.43 |
| numerical | 70.00 | 32.50 | 32.50 | 42.50 | 25.00 | 57.50 | 57.50 | 37.50 | |
| translation | 100.00 | 42.50 | 62.50 | 52.50 | 67.50 | 70.00 | 102.50 | 77.50 | |
| code | 75.00 | 70.00 | 70.00 | 52.50 | 57.50 | 62.50 | 85.00 | 60.00 | |
| MME-OH | 291.67 | 343.33 | 343.33 | 335.00 | 316.67 | 325.00 | 351.67 | 323.33 | |
| MME-AH | 280.00 | 320.00 | 320.00 | 285.00 | 253.33 | 298.33 | 308.33 | 308.33 | |
| Reasoning | 357.14 | 260.00 | 277.86 | 269.64 | 284.29 | 305.71 | 361.43 | 286.43 | |
| LLaVA-v1.5-13B | Default | Temp | Top-k | Top-p | Overall | |
| Object-level Hallucination | existence | 185.0 | 195.0 | 185.0 | 195.0 | 195.0 |
| count | 155.0 | 160.0 | 155.0 | 160.0 | 160.0 | |
| Attribute-level Hallucination | position | 133.3 | 141.7 | 141.7 | 141.7 | 141.7 |
| color | 170.0 | 180.0 | 170.0 | 180.0 | 180.0 | |
| Reasoning | commonsense-reasoning | 127.9 | 136.4 | 125.7 | 135.0 | 136.4 |
| numerical-calculation | 42.5 | 77.5 | 72.5 | 70.0 | 77.5 | |
| text-translation | 77.5 | 125.0 | 92.5 | 105.0 | 125.0 | |
| code-reasoning | 47.5 | 90.0 | 115.0 | 105.0 | 115.0 | |
| All | 938.7 | 1105.6 | 1057.4 | 1091.7 | 1130.6 | |
| LLaVA-v1.5-7B | Default | Temp | Top-k | Top-p | Overall | |
| Object-level Hallucination | existence | 175.0 | 195.0 | 190.0 | 195.0 | 195.0 |
| count | 138.3 | 168.3 | 155.0 | 160.0 | 168.3 | |
| Attribute-level Hallucination | position | 128.3 | 143.3 | 131.7 | 150.0 | 150.0 |
| color | 131.7 | 180.0 | 170.0 | 175.0 | 180.0 | |
| Reasoning | commonsense-reasoning | 102.9 | 122.1 | 120.7 | 113.6 | 122.1 |
| numerical-calculation | 60.0 | 77.5 | 90.0 | 95.0 | 95.0 | |
| text-translation | 92.5 | 107.5 | 120.0 | 120.0 | 120.0 | |
| code-reasoning | 62.5 | 102.5 | 112.5 | 95.0 | 112.5 | |
| All | 891.2 | 1096.3 | 1089.9 | 1103.6 | 1143.0 | |

### v2

@@ 3.3 Other Findings and Analysis
Navigating Challenges in Debiasing LLM-Biased Benchmarks: As illustrated in Fig.˜10, we unravel the intricate interplay between debiasing methods and benchmarks influenced by LLM biases, with Naive’s performance centered around textual information. The post-hoc debiasing methods proposed exhibit inconsistent enhancements on the MMMU benchmark. Upon meticulous examination, when we entirely remove vision information by replacing the input image with an empty string (None∗) or the <unk> identifier (Unk∗), the model maintains comparable or superior performance in specific subsets. In these scenarios, our post-hoc methods, crafted to alleviate biases in LVLMs, fall short of yielding optimal results. Especially in instances where pure LLMs outshine, the application of post-hoc methods may prove detrimental. It is crucial to emphasize that benchmarks tailored for LVLMs should prioritize reliance on input images rather than solely on textual content. Consequently, our methods can act as valuable indicators; when the proposed debiasing methods exhibit suboptimal performance, it signals that the benchmark may tilt more towards LLMs and may not be apt for evaluating LVLMs effectively. In such cases, the proposed sampling methods, VDD, also demonstrate inferiority compared to Naive, with an average accuracy of versus .
@@ 0.C.1 Qualitative examples
We present examples in Fig.˜15 to illustrate that the default decoding strategy leads to significant hallucinations in some instances, resulting in a low GPT-4 evaluation score (2). In these cases, VDD proves highly effective in alleviating hallucinations and achieving higher scores. A similar pattern is observed in Fig.˜16, indicating that merely scaling up the model cannot completely eliminate unrealistic patterns, while VDD consistently delivers superior performance. Examples in Fig.˜17 further affirm the efficacy and accuracy of VDD, generating more precise illustrations.
@@ Appendix 0.D Limitations and Future Work
| LLaVA-v1.5-7B | |||||||||||
| Subsets | Naive | Post-Hoc Debias | Debias Sampling | ||||||||
| None | Unk | Both | VCD[21] | VDD-None | VDD-UNK | Both | |||||
existence | 175.00 | 190.00 | 190.00 | 183.33 | 185.00 | 185.00 | 185.00 | 190.00 | ||
| count | 138.33 | 145.00 | 138.33 | 120.00 | 118.33 | 131.67 | 133.33 | 138.33 | |||
position | 128.33 | 145.00 | 140.00 | 130.00 | 128.33 | 131.67 | 133.33 | 126.67 | ||
| color | 131.67 | 170.00 | 165.00 | 145.00 | 155.00 | 165.00 | 158.33 | 165.00 | |||
| Reasoning | commonsense | 102.86 | 95.00 | 94.29 | 105.00 | 102.14 | 92.14 | 120.00 | 113.57 | ||
| numerical | 60.00 | 70.00 | 67.50 | 45.00 | 92.50 | 70.00 | 77.50 | 70.00 | |||
| translation | 92.50 | 50.00 | 50.00 | 47.50 | 70.00 | 72.50 | 92.50 | 75.00 | |||
| code | 62.50 | 87.50 | 87.50 | 85.00 | 47.50 | 45.00 | 47.50 | 87.50 | |||
| MME-OH | 313.33 | 335.00 | 328.33 | 303.33 | 303.33 | 316.67 | 318.33 | 328.33 | |||
| MME-AH | 260.00 | 315.00 | 305.00 | 275.00 | 283.33 | 296.67 | 291.67 | 291.67 | |||
| Reasoning | 317.86 | 302.50 | 299.29 | 282.50 | 312.14 | 279.64 | 337.50 | 346.07 | |||
| LLaVA-v1.5-13B | |||||||||||
| Subsets | Naive | Post-Hoc Debias | Debias Sampling | ||||||||
| None | Unk | Both | VCD[21] | VDD-None | VDD-UNK | Both | |||||
existence | 180.00 | 188.33 | 188.33 | 190.00 | 173.33 | 190.00 | 188.33 | 190.00 | ||
| 111.67 | 155.00 | 155.00 | 145.00 | 143.33 | 135.00 | 163.33 | 133.33 | ||||
position | 130.00 | 165.00 | 165.00 | 155.00 | 123.33 | 138.33 | 143.33 | 143.33 | ||
| 150.00 | 155.00 | 155.00 | 130.00 | 130.00 | 160.00 | 165.00 | 165.00 | ||||
| Reasoning | commonsense | 112.14 | 115.00 | 112.86 | 122.14 | 134.29 | 115.71 | 116.43 | 111.43 | ||
| numerical | 70.00 | 32.50 | 32.50 | 42.50 | 25.00 | 57.50 | 57.50 | 37.50 | |||
| translation | 100.00 | 42.50 | 62.50 | 52.50 | 67.50 | 70.00 | 102.50 | 77.50 | |||
| code | 75.00 | 70.00 | 70.00 | 52.50 | 57.50 | 62.50 | 85.00 | 60.00 | |||
| MME-OH | 291.67 | 343.33 | 343.33 | 335.00 | 316.67 | 325.00 | 351.67 | 323.33 | |||
| MME-AH | 280.00 | 320.00 | 320.00 | 285.00 | 253.33 | 298.33 | 308.33 | 308.33 | |||
| Reasoning | 357.14 | 260.00 | 277.86 | 269.64 | 284.29 | 305.71 | 361.43 | 286.43 | |||
| LLaVA-v1.5-13B | Default | Temp | Top-k | Top-p | Overall | |||
existence | 185.0 | 195.0 | 185.0 | 195.0 | 195.0 | ||
| count | 155.0 | 160.0 | 155.0 | 160.0 | 160.0 | |||
position | 133.3 | 141.7 | 141.7 | 141.7 | 141.7 | ||
| color | 170.0 | 180.0 | 170.0 | 180.0 | 180.0 | |||
| Reasoning | commonsense-reasoning | 127.9 | 136.4 | 125.7 | 135.0 | 136.4 | ||
| numerical-calculation | 42.5 | 77.5 | 72.5 | 70.0 | 77.5 | |||
| text-translation | 77.5 | 125.0 | 92.5 | 105.0 | 125.0 | |||
| code-reasoning | 47.5 | 90.0 | 115.0 | 105.0 | 115.0 | |||
| All | 938.7 | 1105.6 | 1057.4 | 1091.7 | 1130.6 | |||
| LLaVA-v1.5-7B | Default | Temp | Top-k | Top-p | Overall | |||
existence | 175.0 | 195.0 | 190.0 | 195.0 | 195.0 | ||
| count | 138.3 | 168.3 | 155.0 | 160.0 | 168.3 | |||
position | 128.3 | 143.3 | 131.7 | 150.0 | 150.0 | ||
| color | 131.7 | 180.0 | 170.0 | 175.0 | 180.0 | |||
| Reasoning | commonsense-reasoning | 102.9 | 122.1 | 120.7 | 113.6 | 122.1 | ||
| numerical-calculation | 60.0 | 77.5 | 90.0 | 95.0 | 95.0 | |||
| text-translation | 92.5 | 107.5 | 120.0 | 120.0 | 120.0 | |||
| code-reasoning | 62.5 | 102.5 | 112.5 | 95.0 | 112.5 | |||
| All | 891.2 | 1096.3 | 1089.9 | 1103.6 | 1143.0 | |||